OpenStudy (anonymous):

You are assigned a 4 digit personal identification number(pin) for use with your bank card. Each of the digits can be a whole number from 0 to 9, and the digits can repeat. What is the probability that your Pin is 1234?

OpenStudy (anonymous):

@SmoothMath help please, im not sure how to start this off :o

OpenStudy (anonymous):

1/9* 1/10 1/10 1/10 =1/9000

OpenStudy (anonymous):

4 events. The event that the first digit is 1. The event that the second digit is 2. The event that the third digit is 3. The event that the fourth digit is 4.

OpenStudy (anonymous):

so i have to multiply all 4 together to find out the answer?

OpenStudy (anonymous):

Mhmm =) But first, what is the probability for each of those events?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/10 for each of them?

OpenStudy (anonymous):

prob of evnt 1 =1/9 since 1st digit cant be zero nd prob of remaining other events =1/10

OpenStudy (anonymous):

Wrong, Saura, and I'd also say to you that giving someone the answer right away is a really poor way of helping them.

OpenStudy (anonymous):

1/10 for each. The problem states that each of the digits can be any number from 0 to 9.

OpenStudy (anonymous):

if the 1st digit becomes 0 then it will be a 3 digit no.

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

Bad tutor is bad. Ignore him.

OpenStudy (anonymous):

ok so 1/10 * 1/10 * 1/10 * 1/10 = 1/1000 the probability of choosing 1234?

OpenStudy (anonymous):

lol i will

OpenStudy (anonymous):

Correct =)

OpenStudy (anonymous):

that's it? 1/1000?

OpenStudy (anonymous):

Oops, you made a small mistake. (1/10)*(1/10)*(1/10)*(1/10) = (1/10000)

OpenStudy (anonymous):

oops yes that was my fault i left out a 0

OpenStudy (anonymous):

thanks once again you explain things realllyyy good and just don't give out the answer which is not really learning how to do it. i just have one more problem :s i'll try to solve it on my own and then ask for help if i dont get it

OpenStudy (anonymous):

Sounds good =D